# Improving Algorithmic Efficiency using Cryptography

### v1

@@ Abstract
As a concrete demonstration, we construct a distribution of trapdoored matrices with the following properties: (a) computationally bounded adversaries cannot distinguish a random matrix from one drawn from this distribution, and (b) given a secret key, we can multiply such a $n\times n$ matrix with any vector in near-linear (in $n$) time. We provide constructions both over finite fields and the reals. This enables a broad speedup technique: any algorithm relying on a random matrix—such as those using various notions of dimensionality reduction—can replace it with a matrix from our distribution, achieving computational speedups while preserving correctness.
@@ 1 Introduction
Cryptographic techniques have long played a foundational role in securing communication, ensuring privacy, and enabling secure computation. However, beyond their traditional cryptographic applications, such techniques have also been leveraged to achieve non-cryptographic objectives. Notable examples include using cryptographic hardness assumptions or notions inspired by cryptography to reduce randomness in algorithms (e.g., [BM82, HILL99, NW94, IW97]) or eliminate interaction in protocols (e.g., the celebrated Fiat-Shamir heuristic [FS86]). These applications illustrate how computational hardness assumptions can provide algorithmic benefits beyond security.
@@
To illustrate this approach, we introduce the notion of trapdoored matrices, a class of pseudorandom matrices that allow efficient multiplication when a secret key is available. We construct a distribution of such matrices that satisfies the following key properties:
  • Indistinguishability: A computationally bounded adversary cannot distinguish between a truly random matrix and one sampled from our distribution.
  • Efficient Multiplication: Given a secret key, matrix-vector multiplication can be performed in near-linear time.
We provide constructions of trapdoored matrices over both finite fields and the reals. Over a finite field, our construction is based on the hardness of the Learning Parity with Noise problem, a common cryptographic assumption. Over the reals, we introduce and motivate a hardness assumption based on Kac’s random walks.
As a consequence, any algorithm that relies on a random matrix—such as those used in dimensionality reduction or sketching—can instead use a Trapdoored Matrix, achieving significant speedups due to its efficient multiplication. Crucially, the cryptographic notion of indistinguishability enables a black-box reduction from the correctness of the modified algorithm to that of the original. In particular, the reduction is independent of the specific properties of the random matrix required by the algorithm and its proof.
@@ 1.1 Applications of Random Matrices in Algorithm Design
Random matrices play a fundamental role in modern algorithm design, particularly in settings where dimensionality reduction, efficient approximation, or sketching techniques are required. One of the most well-known applications is the Johnson-Lindenstrauss (JL) Lemma, which states that a set of $n$ points in a possibly high-dimensional Euclidean space can be embedded into a $\Theta\left(\log n/\varepsilon^{2}\right)$-dimensional space while approximately preserving pairwise Euclidean distances up to $\left(1\pm\varepsilon\right)$ multiplicative error. The standard JL transform relies on mapping the set of points through a random projection matrix, often drawn from Gaussian or sub-Gaussian distributions. [JL84]
Beyond the JL lemma, random matrices and the idea of randomized dimensionality reductions are used in various other algorithmic settings, including:
  • Sketching and Streaming: Randomized sketching techniques, such as AMS, CountSketch, and various other streaming algorithms, enable efficient approximation of high-dimensional data with lower memory and computation costs. [AMS96, CCFC02]
@@
  • Subspace Embeddings: Random matrices can be used to construct oblivious subspace embeddings [Sar06, W+14], which facilitate efficient least squares regression, low-rank approximation, and optimization problems by approximately preserving the geometry of a subspace.
@@
  • Error-Correcting Codes (ECCs): Random matrices play a fundamental role in coding theory, particularly in the construction of random linear codes, where a randomly generated generator matrix is used to encode messages into redundant codewords for error correction. [Var57]
  • Dimensionality Reduction Beyond JL: The JL lemma is optimal in the sense that $\Omega\left(\log n/\varepsilon^{2}\right)$ dimensions are needed to preserve all pair-wise distances for any Euclidean set of $n$ points. [LN17] On the other hand, if additional assumptions are made regarding the set of points, a reduction to a lower dimension is sometimes possible. For example, better bounds exist when the set of points is assumed to be of low intrinsic dimension (e.g., doubling dimension) or when not all pair-wise distances are required to be preserved. [IN07, BZD10, CEM+15, MMR19, NSIZ21]
@@
To mitigate this bottleneck, researchers have sought alternative constructions that preserve the key properties of random matrices while allowing for more efficient computation. One prominent example is the extensive study of Fast JL Transforms (e.g., [AC09, AL13, BK21, JPS+22]), which provide fast, low-dimensional embeddings that preserve pairwise distances, similar to traditional Johnson-Lindenstrauss (JL) random projections, but with significantly reduced computational cost—culminating in constructions that run in near-linear time. Another notable example is the development of error-correcting codes with fast encoding.
@@
However, these tailored constructions lack the full flexibility of truly random linear transformations. A fast transform designed to preserve pairwise distances may not necessarily maintain other structural properties, such as spectral norms or subspaces, making it unsuitable for the applications it was not specifically designed for. Our core idea is to replace these randomly sampled matrices with structured pseudorandom matrices—trapdoored matrices—that retain the algorithmic utility of fully random matrices while allowing for faster multiplication. This enables a black-box transformation: any algorithm that relies on random matrices can be accelerated without requiring new, problem-specific correctness proofs, making our approach broadly applicable across a wide range of applications.
1.2 A Specific Application: Faster Classification Inference
We highlight an application where existing fast transforms are not known, demonstrating the versatility of our approach. It was recently shown that for the task of preserving the optimal clustering of a point set, a random projection to a much lower dimension than the JL bound suffices. Optimal bounds on the dimension were proven when parametrized by the number of clusters [MMR19] or by an intrinsic dimension of the point set [NSIZ21]. These results rely on intricate arguments beyond standard applications of the JL lemma.
While this method accelerates training, it does not improve the inference time - the time required to classify an input - since applying a random linear transformation remains as costly as applying an arbitrary one. By replacing the random transformation with a trapdoored matrix, we not only retain the training-time savings but also achieve fast inference by reducing the cost of applying the initial transformation.
Note:
In a concurrent and independent work, Braverman and Newman [^1] A reference to their work will be added in subsequent versions after both papers will become publicly available. suggest and use a construction very similar to that of our Section 3, in a different context motivated by privacy-preserving delegation of computation.
2 Definitions and Results
@@
We provide several constructions of trapdoored matrices over different fields and based on different cryptographic assumptions. We begin with a rigorous definition of trapdoored matrices.
Definition 2.1 (Trapdoored Matrices over $\mathbb{F}^{n\times m}$).
To be considered good, we want the distribution to have efficient multiplication and indistinguishability, which we define next.
Definition 2.2 (Efficient Multiplication).
Definition 2.3 (Indistinguishability).
A Trapdoored Matrix distribution $\mathcal{D}$ is indistinguishable from the uniform distribution $U$ if every polynomial-time algorithm $\mathcal{A}$ is unable to distinguish whether it is given sample access to uniformly drawn random matrices from $U$ or to matrices $M$ drawn using $\mathcal{D}$, that is
$\left|\Pr\left[\mathcal{A}(U)=1\right]-\Pr\left[\mathcal{A}(\mathcal{D})=1\right]\right|\leq\mu(n),$
@@
The main conceptual contribution of our work is the following simple yet powerful insight: computational indistinguishability ensures that any algorithm using random matrices can instead use matrices drawn from a trapdoored matrix distribution without affecting correctness or success probability. Any observable difference in performance would itself serve as a polynomial-time distinguisher between the uniform and trapdoored matrix distributions, contradicting indistinguishability.
Thus, we can replace matrix multiplications within the algorithm with the corresponding trapdoor circuits while preserving correctness, as these circuits produce exactly the same outputs – just more efficiently.
@@
We also note that while we crucially rely on the full generality of indistinguishability to preserve correctness in a black-box manner, our reliance on hardness assumptions is actually weaker than in most works in classical cryptography. Typically, in cryptographic settings, a computationally bounded adversary is assumed to actively attempt to break the assumption. In contrast, in our work, the potential distinguisher is merely a natural algorithm designed for an unrelated, non-cryptographic problem. As a result, it is even less likely that such a distinguisher exists for any of the hardness assumptions we introduce next.
2.1 Cryptographic Hardness Assumptions
We present three (incomparable) hardness assumptions, each of our constructions depends on one of them.
The Learning Parity with Noise Assumption:
The learning parity with noise (LPN) assumption says that the two distributions above are computationally indistinguishable.
Conjecture 2.1 (LPN).
Note that the security (e.g. the runtime and the distinguishing advantage of $A$) is parameterized by the LPN dimension $k$. The best algorithms known for LPN in dimension $k$ with $2^{O(k/\log k)}$ samples run in time $2^{O(k/\log k)}$ [BKW03]; and with $\mathsf{poly}(k)$ samples runs in time $2^{O(k/\log\log k)}$ [Lyu05], so Conjecture 2.1 is on the safer side. The more aggressive subexponential LPN assumption says the following:
Conjecture 2.2 (Subexponential LPN).
The McEliece Assumption Family:
where $\mathsf{Perm}(n)$ is the set of all $n\times n$ permutation matrices.
For a reader familiar with the McEliece encryption scheme, we remark that the assumption as formulated above is sufficent for the security of the McEliece cryptosystem (in conjunction with the LPN assumption over $\mathbb{F}$), but is not necessary for it.
McEliece formulated the assumption with $\mathcal{C}$ being the Goppa code [McE78]; this still remains unbroken and is the basis of candidate encryption schemes submitted to the NIST post-quantum cryptography standardization process [Cla23]. The McEliece assumption w.r.t. “quasi-cyclic” codes have been considered in the literature, is the basis of two different NIST encryption candidates [bik17, hqc17]. All these choices of codes permit near-linear time encoding: that is, given a vector $\mathbf{v}\in\mathbb{F}^{k}$, one can compute $\mathbf{G}\mathbf{v}\in\mathbb{F}^{n}$ in time $\tilde{O}(n+k)$, as opposed to the trivial $O(nk)$. On the other hand, when $\mathcal{C}$ is either the Reed-Solomon or the Reed-Muller code, the assumption is broken via an attack of Sidelnikov and Shestakov [SS92].
Our assumption will be that the McEliece assumption holds w.r.t. Goppa, quasi-cyclic or Hamming quasi-cyclic codes [Cla23, bik17, hqc17]. As in the case of the LPN assumption, we can assume either the (conservative) polynomial or the (aggressive) subexponential versions of the assumption, with different results.
Psuedorandomness of Kac’s Random Walks:
Let $Q_{0}=I_{n}\in\mathbb{R}^{n\times n}$ be the $n$-dimensional identity matrix. For any pair of distinct indices $i,j\in[n]$ and any angle $\theta\in[0,2\pi)$ we denote by $R_{i,j,\theta}$ the rotation of the plane spanned by the basis vectors $e_{i},e_{j}$ by angle $\theta$, that is, the sole linear transformation for which
$R_{i,j,\theta}(e_{k})=e_{k}$ $\forall k\notin\{i,j\}$
$R_{i,j,\theta}(e_{i})=\cos\theta\cdot e_{i}+\sin\theta\cdot e_{j}\;$
$R_{i,j,\theta}(e_{j})=-\sin\theta\cdot e_{i}+\cos\theta\cdot e_{j}\;.$
We define a random walk starting from $Q_{0}=I_{n}$ in which every step consists of a multiplication with a matrix $R_{i,j,\theta}$ where the indices $i,j$ and the angle $\theta$ are drawn uniformly. That is, $i_{t},j_{t}\leftarrow[n]$, $\theta_{t}\leftarrow[0,2\pi)$ and $Q_{t}=R_{i_{t},j_{t},\theta_{t}}$. The resulting matrix is the state $Q_{T}$ of the random walk after $T$ steps.
Conjecture 2.3.
There exists some $T=n\;\emph{poly}(\log n)$ such that the distribution of $Q_{T}$ is computationally indistinguishable from the distribution $U$ in which every column is an i.i.d. unit vector in $S^{n-1}$.
2.2 Our Results
@@
Our third construction is for the distribution $U$ of real-valued matrices in which each column is a random unit vector of $S^{n-1}$. This is the distribution for which Johnson-Lindenstrauss (JL) like dimension reduction algorithms work.
@@
Under the psuedorandomness of Kac’s random walks, there is a collection of trapdoored matrices over the distribution $U$ for which, given the trapdoor, $n\times n$ matrix-vector multiplication can be performed in time $n\cdot\mathsf{poly}(\log n)$.
@@
3 Trapdoored Matrices from LPN

### v2

@@ Abstract
As a concrete demonstration, we construct a distribution of trapdoored matrices with the following properties: (a) computationally bounded adversaries cannot distinguish a random matrix from one drawn from this distribution (under computational hardness assumptions), and (b) given a trapdoor, we can multiply such an $n\times n$ matrix with any vector in near-linear (in $n$) time. We provide constructions both over finite fields and over the reals. This enables a broad speedup technique: any algorithm relying on a random matrix—such as those that use various notions of dimensionality reduction—can replace it with a matrix from our distribution, achieving computational speedups while preserving correctness.
Using these trapdoored matrices, we present the first uniform reduction from worst-case to approximate and average-case matrix multiplication with optimal parameters (improving on Hirahara-Shimizu STOC 2025, albeit under computational assumptions), the first worst-case to average-case reductions for matrix inversion, solving a linear system, and computing a determinant, as well as a speedup of inference time in classification models.
@@ 1 Introduction
Cryptographic techniques have long played a foundational role in securing communication, ensuring privacy, and enabling secure computation. However, beyond their traditional cryptographic applications, such techniques have also been leveraged to achieve non-cryptographic objectives. Notable examples include using cryptographic hardness assumptions or notions inspired by cryptography to reduce randomness in algorithms (e.g. pseudorandom generators starting from Blum-Micali [BM82, HILL99, NW94, IW97]), and eliminate interaction in protocols (e.g. the celebrated Fiat-Shamir heuristic [FS86, CCH+19]). These applications illustrate how computational hardness assumptions can provide algorithmic benefits beyond the obvious goals of security and privacy.
@@
To illustrate this approach, we introduce the notion of trapdoored matrices, a class of pseudorandom matrices that allow efficient multiplication when a trapdoor is available. We construct a distribution of such $n\times n$ matrices that satisfies the following key properties:
  • Indistinguishability: A computationally bounded adversary cannot distinguish between a truly random matrix and one sampled from our distribution; and
  • Efficient Multiplication: Given a trapdoor, matrix-vector multiplication can be performed in near-linear $\tilde{O}(n)$ time. As an immediate consequence, matrix-matrix multiplication can be performed in near-quadratic $\tilde{O}(n^{2})$ time.
We provide constructions of trapdoored matrices over both finite fields and over the reals. Over a finite field, we provide two constructions: one whose indistinguishability holds under the hardness of the well-known learning parity with noise problem (LPN) [BFKL93]; and a second based on the McEliece assumption, the basis of McEliece’s cryptosystem from the 1970s [McE78]. Over the reals, our construction relies on the pseudorandomness of the Kac random walk [Kac56]; we introduce and motivate the relevant hardness assumption.
As a consequence, any algorithm that relies on a random matrix—such as those used in dimensionality reduction or sketching—can instead use a trapdoored matrix, achieving significant speedups due to the availability of efficient multiplication. Crucially, the cryptographic notion of indistinguishability enables a black-box reduction from the correctness of the modified algorithm to that of the original. In particular, the reduction is independent of the specific properties of the random matrix required by the algorithm and its correctness proof. We remark that in fact, the hardness assumptions required for algorithmic correctness are milder than what one typically assumes in cryptography, e.g. assuming that LPN is hard for cubic-time (as opposed to any polynomial-time) algorithms suffices if the underlying algorithm runs in cubic time. Moreover, we use these hardness assumptions in a conceptually weaker manner than their standard use in cryptography: usually, it is assumed that a computationally-bounded yet possibly highly motivated player tries to break the conjecture. In our case, the pseudorandomness stemming from the cryptographic conjectures is not meant to trick a dedicated adversary, but an innocent algorithm designed to solve a completely non-cryptographic problem.
@@ 1.1 Applications of Random Matrices in Algorithm Design
Random matrices play a fundamental role in modern algorithm design, particularly in settings where dimensionality reduction, efficient approximation, or sketching techniques are required. One of the most well-known applications is the Johnson-Lindenstrauss (JL) Lemma, which states that a set of $N$ points in a possibly high-dimensional Euclidean space can be embedded into a $\Theta\left((\log N)/\varepsilon^{2}\right)$-dimensional space while approximately preserving pairwise Euclidean distances up to $\left(1\pm\varepsilon\right)$ multiplicative error. The standard JL transform relies on mapping the set of points through a random projection matrix, often drawn from Gaussian or sub-Gaussian distributions [JL84, IM98].
Beyond the JL lemma, random matrices and the idea of randomized dimensionality reduction are used in various other algorithmic settings, including:
  • Sketching and Streaming: Randomized sketching techniques, such as AMS, CountSketch, and various other streaming algorithms, enable efficient approximation of high-dimensional data with lower memory and computation costs [AMS96, CCFC02].
@@
  • Subspace Embeddings: Random matrices are used to construct oblivious subspace embeddings [Sar06, W+14], which facilitate efficient least squares regression, low-rank approximation, and optimization problems by approximately preserving the geometry of a subspace.
@@
  • Error-Correcting Codes (ECCs): Random matrices play a fundamental role in coding theory, particularly in the construction of random linear codes, where a randomly generated generator matrix is used to encode messages into redundant codewords for error correction [Var57].
  • Dimensionality Reduction Beyond JL: The JL lemma is optimal in the sense that $\Omega\left((\log N)/\varepsilon^{2}\right)$ dimensions are needed to preserve all pair-wise distances for any Euclidean set of $N$ points [LN17]. On the other hand, if additional assumptions are made regarding the set of points, a reduction to a lower dimension is sometimes possible. For example, better bounds exist when the set of points is assumed to be of low intrinsic dimension (e.g., doubling dimension) or when not all pair-wise distances are required to be preserved [IN07, BZD10, CEM+15, MMR19, NSIZ21].
@@
To mitigate this bottleneck, researchers have sought alternative constructions that preserve the key properties of random matrices while allowing for more efficient computation. One prominent example is the extensive study of fast JL transforms (e.g., [AC09, AL13, BK21, JPS+22]), which provide fast, low-dimensional embeddings that preserve pairwise distances, similar to traditional Johnson-Lindenstrauss (JL) random projections, but with significantly reduced computational cost—culminating in constructions that run in near-linear time. Another notable example is the development of error-correcting codes with fast encoding [GDP73, Spi95].
@@
However, these tailored constructions lack the full flexibility of truly random linear transformations. A fast transform designed to preserve pairwise distances may not necessarily maintain other structural properties, such as spectral norms or subspaces, making it unsuitable for the applications it was not specifically designed for. Our core idea is to replace these randomly sampled matrices with structured, yet pseudorandom matrices—trapdoored matrices—that retain the algorithmic utility of fully random matrices while allowing for faster multiplication. This enables a black-box transformation: any algorithm that relies on random matrices can be accelerated without requiring new, problem-specific correctness proofs, making our approach broadly applicable across a wide range of applications.
1.2 Our Results
@@
We now present our results in more detail, starting with our notion of trapdoored matrices. We then provide several constructions of trapdoored matrices over different fields and based on different cryptographic assumptions.
An efficiently samplable distribution $\mathcal{D}$ over pairs $(M,C)$ where $M$ is a matrix over $\mathbb{F}^{n\times m}$ (where $\mathbb{F}$ is a field, either a finite field or the reals in this work) and $C:\mathbb{F}^{m}\to\mathbb{F}^{n}$ is a circuit, is a trapdoored matrix distribution if:
  • (Correctness)  For every vector $v\in\mathbb{F}^{n}$, the circuit $C$ computes the matrix-vector multiplication
    $C(v)=Mv.$
  • (Efficiency)  The circuit $C$ is small: it has size $\widetilde{O}(n+m)$.
  • (Indistinguishability)  A matrix $M$ drawn from $\mathcal{D}$ is computationally indistinguishable from a uniformly random matrix over $\mathbb{F}^{n\times m}$ (or from an appropriate finite distribution, e.g. a truncated and discretized Gaussian, if the field is the reals).
In other words, the circuit $C$ is the trapdoor (associated to $M$) that allows for quick computation of the matrix-vector product with any given vector $v$.
@@
The main conceptual contribution of our work is the following simple yet powerful insight: computational indistinguishability ensures that any algorithm using random matrices can instead use matrices drawn from a trapdoored matrix distribution without affecting correctness or success probability. Any observable difference in performance would itself serve as a polynomial-time distinguisher between the uniform and trapdoored matrix distributions, contradicting indistinguishability. Thus, we can replace matrix multiplications within the algorithm with the corresponding trapdoor circuits while preserving correctness, as these circuits produce exactly the same outputs – just more efficiently.
@@
We also note that while we crucially rely on the full generality of indistinguishability to preserve correctness in a black-box manner, our reliance on hardness assumptions is actually weaker than in most works in classical cryptography. Typically, in cryptographic settings, a computationally bounded adversary is assumed to actively attempt to break the assumption. In contrast, in our work, the potential distinguisher is merely a natural algorithm designed for an unrelated, non-cryptographic problem. As a result, it is even less likely that such a distinguisher exists for any of the hardness assumptions we employ in this paper.
Our Constructions of Trapdoored Matrices.
@@
The learning parity with noise (LPN) assumption is a classical average-case hardness assumption originating from learning theory and coding theory [BFKL93, Ale03]. For a detailed discussion of the assumption, we refer the reader to Section 3.1.
@@
The McEliece assumption is a formalization of an assumption underlying the security of the McEliece public-key cryptosystem [McE78] from the 1970s, also the basis of the NIST post-quantum cryptography submission “Classic McEliece” [Cla23]. For a detailed discussion of the assumption, we refer the reader to Section 4.1.
Our third construction is for distributions of real-valued matrices, such as those used in JL-like dimension reductions. We say that a distribution $\mathcal{D}$ over matrices is Haar-invariant if the matrix products $O\cdot A$ and $A\cdot O$, where $A\sim\mathcal{D}$ and $O$ is sampled from the Haar-measure over $SO(n)$, are both distributed as $\mathcal{D}$. Informally, these are all distributions that are invariant to a random rotation of either the input or output space. An example for such a distribution $\mathcal{D}$ is the distribution of matrices in which every coordinate is an i.i.d. normal Gaussian variable. We say that a distribution $\mathcal{D}$ over symmetric matrices is symmetrically Haar-invariant if the matrix product $O\cdot A\cdot O^{-1}$, where $A\sim\mathcal{D}$ and $O$ is sampled from the Haar-measure over $SO(n)$, is distributed as $\mathcal{D}$. Informally, these are all distributions that are invariant to conjugation by a random rotation of the space. An example of such a distribution is the Gaussian Orthogonal Ensemble (GOE).
@@
Let $\mathcal{D}$ be any distribution that is either Haar-invariant or symmetrically Haar-invariant. Under the psuedorandomness of Kac’s random walks, there is a collection of trapdoored matrices over the distribution $\mathcal{D}$ for which, given the trapdoor, $n\times n$ matrix-vector multiplication can be performed in time $n\cdot\mathsf{poly}(\log n)$.
For a detailed discussion of Kac’s random walks and their plausible pseudorandomness, we refer the reader to Section 5.1. A similar assumption in the context of finite fields was studied by Sotiraki [Sot16] in the context of constructing fine-grained key-exchange protocols in the spirit of Merkle [Mer78].
@@
Beyond this starting point, our constructions require further ideas. For example, our basic LPN-based construction achieves $n^{3/2}$-time matrix-vector multiplication, worse than our eventual goal. We augment this basic construction with recursion to achieve near-linear-time computation. The same goes for our construction based on the McEliece assumption. Moving on to trapdoor matrices over the reals, our basic and natural construction samples trapdoored matrices for the uniform distribution over $SO(n)$, the special orthogonal group. We then augment this basic construction with a reduction showing how to use such a sampler to construct trapdoored matrices for any Haar-invariant distribution, in particular the Gaussian distribution.
1.3 Motivating Applications
1.3.1 Faster Inference for Classification Models
We highlight an application where fast transforms are not known, demonstrating the versatility of our approach. It was recently shown that for the task of preserving the optimal clustering of a point set, a random projection to a much lower dimension than the JL bound suffices. Optimal bounds on the dimension were proven when parametrized by the number of clusters [MMR19] or by an intrinsic dimension of the point set [NSIZ21]. These results rely on intricate arguments beyond standard applications of the JL lemma.
While this method accelerates training, it does not improve the inference time since the cost of applying the random linear transformation for dimensionality reduction is significant in comparison to the cost of running the neural network inference on the lower-dimensional data. By replacing the random linear transformation with a trapdoored matrix, we not only retain the training-time savings but also achieve fast inference by reducing the cost of applying the initial transformation. (This could also reduce the training time, but there, the cost of the initial dimension-reduction of the data— a rectangular matrix multiplication— is often not the dominant cost.)
1.3.2 Uniform Error-Correction for Matrix Multiplication
Given a black-box algorithm that approximately computes matrix multiplication for an average input, can we design a nearly-as-efficient algorithm that computes matrix multiplication exactly for any input? Several variants of this question were extensively studied. Blum, Luby, and Rubinfeld [BLR90] initially showed that given an algorithm that computes $A\cdot B$ exactly for a sufficiently large constant fraction of input pairs $A,B\in\mathbb{F}_{p}^{n\times n}$, then we can also compute $A\cdot B$ for any input pair with high probability using almost the same amount of time. Subsequent works [AGGS22, HS23] obtained the same result even when the given algorithm produces the correct multiplication only on a very small fraction of input pairs.
Later works [GSS24, HS25] the above was generalized to the case where the given algorithm only produces a matrix that is close to $A\cdot B$ in Hamming distance, but is not guaranteed to exactly compute the multiplication for any input pair. Recently, Hirahara and Shimizu [HS25] obtained a non-uniform such reduction given the weakest possible multiplication algorithm over a field of prime order: Given an algorithm that on a random pair $A,B\sim\mathbb{F}_{p}^{n\times n}$ of inputs returns a matrix whose relative Hamming distance from $A\cdot B$ is at most $1-\frac{1}{p}-\varepsilon$, for some $\varepsilon>0$, they can compute $A\cdot B$ exactly for any input pair $A,B$ in a similar time. Note that an algorithm that simply returns a random matrix is in expectation only $(1-\frac{1}{p})$-far from the correct multiplication, and they assume only a small advantage over that. On the other hand, their reduction is non-uniform, meaning that they need some large pre-processing time that depends only on the matrix size $n$. Furthermore, all of the aforementioned results are based on highly sophisticated tools.
Using our construction of trapdoored matrices, we are able to obtain the same result using a uniform reduction, affirmatively answering the main remaining problem in this line of work — under a computational assumption. Furthermore, given a trapdoored matrix family, our reduction is extremely simple and works in any finite field.
Theorem (6.2).
Let $\mathbb{F}_{p}$ be a finite field of order $p$, and assume the existence of a trapdoored matrix family over $\mathbb{F}_{p}$ and let $\varepsilon>0$. Let $\mathcal{M}:\mathbb{F}_{p}^{n\times n}\times\mathbb{F}_{p}^{n\times n}\rightarrow\mathbb{F}_{p}^{n\times n}$ be an algorithm running in time $T(n)$ such that for every sufficiently large $n$,
$\mathbb{E}_{A,B\sim\mathbb{F}_{p}^{n\times n}}\Big{[}\text{dist}\left(A\cdot B,\ \mathcal{M}\left(A,B\right)\right)\Big{]}\leq\left(1-\frac{1}{p}-\varepsilon\right)n^{2}.$
Then, we can construct an algorithm that for any matrices $A,B\in\mathbb{F}_{p}^{n\times n}$ computes the multiplication $A\cdot B$ exactly with high probability, in time $\tilde{O}\left(T(n)\right)$.
1.3.3 Worst Case to Average Case Reductions for Other Linear Algebraic Operations
Similar questions to those discussed above can be asked for other linear algebraic operations beyond matrix multiplication. In Strassen’s seminal paper [Str69], which introduced the first improvement to the matrix multiplication exponent, he also showed that matrix inversion, determinant computation, and Gaussian elimination have the same computational complexity as matrix multiplication (up to lower-order terms). To the best of our knowledge, no worst-case to average-case reductions are known for these problems, even when the average-case algorithm is guaranteed to produce a correct answer with probability at least $99\%$. In [BLR90], the authors described algorithms that solve these problems in the worst case, assuming access to average-case algorithms for all of these problems (including, in particular, average-case matrix multiplication). In contrast, the simplicity of our reductions arising from trapdoored matrices enables reductions that rely only on the average-case version of the same problem.
In Section 6, we present reductions, using trapdoored matrix families such as those we construct in this paper, from the worst-case versions of matrix multiplication, matrix inversion, solving a system of linear equations with a unique solution and computing a determinant to the average-case version of each problem. As in the worst-case all of these problems are computationally equivalent, this implies that an average-case algorithm for any of them would result in a worst-case algorithm for all of them.
1.4 Related Work
Pseudorandomness has been used in the past to reduce space in algorithmic tasks. A prominent example is algorithmic hash functions: instead of using a fully random function, a succinctly represented pseudorandom function often suffices for algorithmic applications-for instance, $k$-wise independent hash families can efficiently replace fully random functions in hash tables. Similarly, certain classes of expander graphs, such as those constructed as Cayley graphs, provide an explicit and succinct representation that eliminates the need to store a fully random graph in memory while preserving key properties. Motivated by these examples, Goldreich, Goldwasser, and Nussboim [GGN10] initiated a broader line of research exploring how to implement large (pseudo-)random objects in a space-efficient manner, while also maintaining efficient access to these objects (such as applying a pseudo-random function on an input or accessing the adjacency lists of a pseudo-random graph). Their and subsequent works develop such implementations for pseudorandom functions and graphs, including both dense [GGN10] and sparse [NN07] instances.
Note:
In a concurrent and independent work, Braverman and Newman [BN25] suggest and use a construction very similar to that of our Section 3, in a different context motivated by privacy-preserving delegation of computation. Our focus, on the other hand, is using these cryptographic primitives for non-cryptographic objectives. Moreover, we present additional constructions based on McEliece-type assumptions and for real-valued matrices.
2 Trapdoored Matrices
We start with a rigorous definition of our notion of trapdoored matrices. We then provide several constructions of trapdoored matrices over different fields and based on different cryptographic assumptions. Let $\mathbb{F}$ be a field (either a finite field or the reals, in this work).
Definition 2.1 (Trapdoored Matrices over $\mathbb{F}^{n\times m}$).
To be considered good, we want the distribution $\mathcal{D}$ to have efficient multiplication and indistinguishability, which we define next.
Definition 2.2 (Efficient Multiplication).
Definition 2.3 (Indistinguishability).
A trapdoored matrix distribution $\mathcal{D}$ is indistinguishable from the uniform distribution $\mathcal{U}$ if every polynomial-time algorithm $\mathcal{A}$ is unable to distinguish whether it is given a uniformly drawn random matrix from $\mathcal{U}$ or a matrix $M$ drawn using $\mathcal{D}$, that is
$\left|\Pr\left[\mathcal{A}(M)=1:M\sim\mathcal{U}\right]-\Pr\left[\mathcal{A}(M)=1:(M,C)\sim\mathcal{D}\right]\right|\leq\mu(n),$
This definition can be generalized in several ways: first, from matrix-vector multiplication to matrix-matrix multiplication; and secondly, by allowing for efficient multiplication on both the right and the left; and thirdly, by allowing for efficient multiplication by both $M$ and $M^{-1}$. All our constructions can be easily extended to matrix-matrix multplication. Our LPN-based construction supports left and right multiplication. Our construction over the reals supports left and right multiplication, as well as multiplication by $M^{-1}$.
Yet another extension that one could conceive of would be to quickly output a pseudorandom matrix together with a property or class of properties, e.g. $M$ together with its determinant, rank, etc. We do not explore this extension further in this work.
3 Trapdoored Matrices from the Learning Parity with Noise Assumption
3.1 The Learning Parity with Noise (LPN) Assumption
The learning parity with noise (LPN) assumption says that the two distributions above are computationally indistinguishable for any $n=\mathsf{poly}(k)$ samples.
Conjecture 3.1 (LPN).
Note that the security (e.g. the runtime and the distinguishing advantage of $A$) is parameterized by the LPN dimension $k$. The LPN assumption (and its generalization over larger fields) has found a number of applications in cryptography, and seems exponentially hard in a wide range of parameters (see e.g.  [LWYY24] and the references therein). The best algorithms known for LPN in dimension $k$ with $2^{O(k/\log k)}$ samples and $p=O(1)$ run in time $2^{O(k/\log k)}$ [BKW03]; and with $\mathsf{poly}(k)$ samples runs in time $2^{O(k/\log\log k)}$ [Lyu05]. On the other hand, with $p=O((\log k)/k)$, the assumption is trivially broken in $\mathsf{poly}(k)$ time. In between, for a general $p$, the best known attacks run in time $2^{\tilde{O}(H(p)\cdot k)}$.
Conjecture 3.1 is on the safer side for $p\gg(\log k)/k$; the more aggressive subexponential LPN assumption conjectures the following, and is plausible for a polynomial $p=k^{-\beta}$ for any constant $\beta<1$ (with a related $\alpha$).
Conjecture 3.2 (Subexponential LPN).
3.2 Construction from LPN